Twice the smaller of two numbers is one-half of the larger number. The larger is 10 more than 3 times the smaller. Find the numbers.
Answer by fractalier(6550)

Call the smaller, x, and the larger, y. Then we have
2x = (1/2)y or 4x = y and we have
y = 10 + 3x
Substituting we have
4x = 3x + 10
x = 10
y = 40